OpenStudy (lukecrayonz):

Of those numbers whose sum is 70, find the two that have the maximum product. That is, maximize Q=xy where x+y=70

OpenStudy (anonymous):

What class is this for?

OpenStudy (anonymous):

@Lukecrayonz what class is this for?

OpenStudy (dumbcow):

x + y = 70 --> y = 70 - x xy = x(70-x) = 70x - x^2 differentiate and set equal to 0 to maximize 70x-x^2

OpenStudy (lukecrayonz):

@FutureMathProfessor business calc

OpenStudy (lukecrayonz):

@dumbcow derivative of 70x-x^2 is 2x-70, so now what?:O

OpenStudy (lukecrayonz):

set that equal to zero and solve for x?

OpenStudy (dumbcow):

yes

OpenStudy (lukecrayonz):

35

OpenStudy (lukecrayonz):

Oh. So just 35 and 35..

OpenStudy (dumbcow):

yep so max Q = 35^2

OpenStudy (lukecrayonz):

1225